The locus of a point equidistant from two given points a and b is given by

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

locus of point will be perpendicular bisector of these two vector.

let the general point on the line is r

so line will pass through (a+b)/2

and r-(a+b)/2     and (b-a)  will be perpendicular

so {r-(a+b)/2 }.(b-a) =0

 r.(b-a) =1/2( |b|2 -|a|2)

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin